5. You have received the following data involving milling of a part by a worker in your department. The data are expressed in minutes: Work 50 68 30 57 90 70 68 67 80 30 31 30 Your department is scheduled for one 8-hour shift, 5 days per week and workers are given a daily allowance of 72 minutes for personal time and other contingencies. Contractually daily standards are calculated as the normal time divided by 1-56 allowance. Assembly workers are paid $10.50 per hour. Historically scrap on this part has been averaging 8%, A. Calculate the standard time to produce on part based on the above time study information. B. You are told that the schedule calls for 550 units per day in the foreseeable future. How many workers are needed to achieve this schedule?
Answer #1
Combined Average
Work Element Average Seconds
1 0.500 30
2 0.250 15
3 0.680 40.8
4 0.550 33
5 0.300 18
Total 2.280 136.800
Equals 2.28Mins

B) 550Units is excluding Scrap

So therefore 550/0.92(After allowance for scrap)=598units per day

Total Time Contributed
by each worker per
week
8hours*60(min)
less
72(Personal Time)*5Daysof
week=
2040Mins
Daily 408mins
Daily Max units per employee
408/2.28
179units

So therefore to manufacture 598units we need 598/179= 3.34 (more than 3 EMployees )

C)

Total Time Contributed
by each worker per
week
8hours*60(min)
less
72(Personal Time)*5Daysof
week=
2040Mins

So therefore with 2040Mins per week per employee . he can manufacture = 2040/2.28= 895Units

With 110% Efficiency and 4 % Scrap Allowance = 895*110%*0.96=945Units

D)

Particulars
Direct Material cost 2.15
Direct Labour 0.4 2.28*10.5/60
Total 2.55
Fixed cost(120%
of above)
3.06
Total 5.61